OpenStudy (anonymous):

MEDAL .. Consider the sequence 1, 2/3, 3/5, 4/7, 5/9 . a.Describe the pattern formed in the sequence. b. Find the next three terms.

OpenStudy (anonymous):

hint: you can write 1 as 1/1 so your sequence is like \[\frac 11 \quad \frac 23 \quad \frac 35 \quad \frac 47 \quad \frac 59\] what do you think is the pattern?

OpenStudy (anonymous):

i get so lost w/ fractions and thats y im completley lost ... :/

OpenStudy (anonymous):

he numerator increases by 1 for each successive term and the denominator increases by 2..?

OpenStudy (anonymous):

yes that's it^

OpenStudy (anonymous):

so the next three terms: 6/11 7/13 8/15

OpenStudy (anonymous):

ok got it!!!! ty @robinnicole218 !

OpenStudy (anonymous):

no prob, im good at algebra 1, that's what it is right? ask any ques~!